# ViMo: A Generative Visual GUI World Model for App Agent

### v1

@@ Abstract
App agents, which autonomously operate mobile Apps through Graphical User Interfaces (GUIs), have gained significant interest in real-world applications. Yet, they often struggle with long-horizon planning, failing to find the optimal actions for complex tasks with longer steps. To address this, world models are used to predict the next GUI observation based on user actions, enabling more effective agent planning. However, existing world models primarily focus on generating only textual descriptions, lacking essential visual details. To fill this gap, we propose ViMo, the first visual world model designed to generate future App observations as images. For the challenge of generating text in image patches, where even minor pixel errors can distort readability, we decompose GUI generation into graphic and text content generation. We propose a novel data representation, the Symbolic Text Representation (STR) to overlay text content with symbolic placeholders while preserving graphics. With this design, ViMo employs a STR Predictor to predict future GUIs’ graphics and a GUI-text Predictor for generating the corresponding text. Moreover, we deploy ViMo to enhance agent-focused tasks by predicting the outcome of different action options. Experiments show ViMo’s ability to generate visually plausible and functionally effective GUIs that enable App agents to make more informed decisions. Project link: https://ai-agents-2030.github.io/ViMo/.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs)  [^2] By LLMs, we refer to the concept of foundation models that accept various input modalities (e.g., visual language models (VLMs), multimodal LLMs (MLLMs)) while producing textual sequences W. contributors (2024). have unlocked new possibilities for deploying AI agents across diverse fields Li et al. (2023); Gou et al. (2023); Rawles et al. (2024b; b); Wang et al. (2024a). A notable application is the smartphone application (App) agents Rawles et al. (2024a); Wang et al. (2024a), designed to directly interact with Graphical User Interfaces (GUIs) to perform tasks autonomously and efficiently in a mobile operating system.
However, existing agents struggle with making decisions for tasks requiring longer steps Chae et al. (2024). To address this “long-horizon” limitation, an increasing number of studies have introduced a world model, which predicts how GUIs evolve in response to user actions Gu et al. (2024). Yet, these models typically rely on language to describe future observations. These language-based descriptions often fail to capture the intricate visual details, such as the location and colour of GUI elements, necessary for a precise representation Chae et al. (2024). There is a need for a visual GUI world model capable of comprehensively representing observations in visual modality. Addressing this gap serves as the primary motivation for this work.
Figure 1: Visualisation of ViMo in generating next GUIs based on the current GUI observation and a user action.
To build a visual GUI world model capable of generating plausible future GUI observations that are visually consistent with user actions, as shown in Fig. 1, a straightforward approach involves generating each pixel of a GUI using image generations  Rombach et al. (2022). Although these methods demonstrate promising results, such as the GUI graphic generation on the location, style, and colour of GUI elements Wei et al. (2024), or scene-text generation in a style that aligns the visual context Chen et al. (2024b); Zhang et al. (2024c), they still display distortions in the text rendering, particularly for small-sized text where each pixel is critical for accurately identifying and representing the text (see Fig. 2 for an illustration).
@@
We evaluate ViMo in two distinct scenarios to comprehensively demonstrate its effectiveness. First, we conduct a world model ability evaluation, focusing on two critical aspects: (a) GUI quality evaluation, where we assess the generated GUIs using a visual similarity score, instructional accuracy score, and action readiness score to ensure a detailed and holistic evaluation of the generated GUI; (b) GUI-based trajectory synthesis evaluation, where the generated GUI is leveraged as the input to an App agent to generate subsequent actions, with higher-quality trajectories indicating a GUI more aligned with real-world environment. These experiments provide a robust assessment of the model’s ability to generate visually precise and contextually plausible GUIs. Second, we test ViMo in an agent-focused task to evaluate its utility in practical applications. Given a goal and the current observation of an App, an agent determines the optimal actions needed to achieve the goal Wang et al. (2024a). By accurately predicting the next GUI given the current GUI and an action, ViMo enables the agent to better anticipate the action outcomes and make more informed decisions. This evaluation highlights the model’s applicability in real-world scenarios and its potential to enhance decision-making for App agents.
@@
$\bullet$ Extensive experiments show the effectiveness of ViMo in both the world model ability evaluation and agent-focused tasks. Specifically, ViMo achieves on average 29.14% relative improvement in performance over existing world models and enhances the App agent’s step-wise action prediction accuracy with a relative performance gain of 14.07%.
@@ 2.1 App Agent
App agents, powered by LLMs, have made significant strides in automating tasks on mobile Apps Wen et al. (2024b); Chen et al. (2024c); Zhang et al. (2024a; b); Lee et al. (2023); Nguyen et al. (2024). These agents interact with GUIs by emulating human actions such as tapping, typing, and swiping. This enables them to accomplish various user goals, such as searching for products or setting alarms. Approaches in this domain are broadly divided into language-based and multi-modality-based methods. Language-based methods rely on textual description of the App observation and the user goal to generate appropriate actions Wen et al. (2024a); Rawles et al. (2024a); Li et al. (2024), while multi-modality-based methods enhance this capability by incorporating GUIs for a more comprehensive understanding of the interface Rawles et al. (2024a); Wang et al. (2024a); Christianos et al. (2024); Wang et al. (2024b). Despite the promising results, these approaches struggle with long-horizon tasks that require multiple interdependent actions and a deep understanding of dynamic environments Chae et al. (2024). For this challenge, recent works Putta et al. (2024); Koh et al. (2024) introduce the use of real-world emulators to simulate GUI changes resulting from agent actions, enabling App agents to better navigate complex, multi-step scenarios and improve decision-making accuracy. However, the use of real-world emulators faces significant drawbacks, including the high computational costs of setting up emulators for specific mobile systems and the safety risks involved in executing real-world interactions, such as repeatedly sending messages or making purchases. To overcome these, world models have attracted increasing attention as a more efficient alternative Chae et al. (2024); Gu et al. (2024).
@@ 2.2 World Model
By observing the real world, world models can predict how the environment evolves in response to an action LeCun (2022); Ding et al. (2024). For instance, video generation models such as Sora OpenAI (2024) simulate future observations by predicting videos given a video observation and GameNGen Valevski et al. (2024) predicts how a game system will respond to user actions. Notably, the ability to anticipate potential outcomes of actions has proven to be highly beneficial in informing decision-making processes Pascanu et al. (2017); Yang et al. (2024); Schrittwieser et al. (2020); Hafner et al. (2019). Inspired by their success, world models have emerged to predict the next observation on websites Chae et al. (2024); Gu et al. (2024). These models Chae et al. (2024); Liu et al. (2023) typically take a website observation and an action as inputs to generate a textual description of the next observation. While websites provide multiple sources of information, including the actual site and their CSS or HTML source files, mobiles present a more limited context, as only the GUIs are typically accessible. Moreover, such text-only descriptions often lack the precise visual details required for accurately predicting future observations, highlighting the need for a visual world model capable of generating high-fidelity future GUI images.
@@ 2.3 GUI Generation
With the rapid advancements in image generation techniques, such as diffusion models Rombach et al. (2022); Kumari et al. (2023); Cao & Gong (2024), generating realistic images has become increasingly feasible. Building on these developments, previous methods have explored generating GUI directly in pixel space. For instance, layout generation methods generate the location of GUI elements Lu et al. (2023); Zheng et al. (2023); Sobolevsky et al. (2023); Zhao et al. (2019), scene-text generation methods generate text that aligns with the visual context Chen et al. (2024a; b); Zhang et al. (2024c); Zeng et al. (2024), UI-diffuser Wei et al. (2024) fine-tune a stable diffusion model to generate mobile GUIs conditioned on text prompts. For the next GUI generation conditioned on current GUI observation and a user action, it seems straightforward to resort to an image-and-text-conditioned approach Brooks et al. (2023). However, we find that pixel-based image generation struggles with rendering text accurately, as even minor pixel prediction errors can lead to distortions, particularly for small-sized text (see Fig. 2 for examples).
@@ 3 Method
Figure 2: GUIs generated by image generation methods (UI-diffuser Wei et al. (2024), InstructPix2Pix Brooks et al. (2023), Textdiffuser-2 Chen et al. (2024b)). Cropped for space efficiency.
@@ 3.2 ViMo: Generative Visual GUI World Model
To tackle the limitation of existing methods Chen et al. (2024b); Brooks et al. (2023); Wei et al. (2024) in generating visually plausible text for a GUI, as shown in Fig. 2, we propose ViMo, a novel generative visual GUI world model that decouples the graphic and text content generation. As shown in Fig. 3, we first detect and remove all the text in the GUI by overlaying it with a text symbol to create the Symbolic Text Representation (STR). Then a STR predictor is leveraged for determining the STR representation of the next GUI with a pixel-based diffusion process. Finally, a GUI-text predictor is proposed to generate the text content for each symbol using an LLM, followed by a handcrafted design to overplay the text into the STR image to create the next GUI. Their details are specified in the following.
@@ 3.2.1 STR: Symbolic Text Representation
To develop a GUI prediction model that eliminates the need to generate specific text content, we propose the Symbolic Text Representation (STR), where all the text content (actual words) within the GUI image is symbolised (overlayed) with uniform text symbols (placeholders). To be specific, we create an STR representation from a given GUI image with three steps: 1) using an OCR model Shi et al. (2016); Qiao et al. (2020) to detect text within the GUI; 2) masking the detected text by overlaying it with a box filled with white and bordered in black; 3) we leverage an LLM to filter out static text displayed on static GUI elements and preserve it in the image, as it does not involve any semantic evolution or dynamic changes and remains unchanged as part of specific elements such as a keyboard or a clock face. Additionally, we empirically find that predicting this static text with complex spatial patterns poses significant challenges for the LLM.
@@ 3.2.2 STR Predictor
Building on the powerful generative capability of diffusion-based models Rombach et al. (2022), we introduce a STR predictor specifically trained to understand a given STR and a user action, enabling it to generate the corresponding next STR effectively. In particular, we fine-tune a pre-trained stable diffusion model Rombach et al. (2022) to predict the next STR, conditioned on the STR of the current GUI and the user action. Given a STR representation ($\text{STR}_{x_{k}}$) extracted from GUI ($x_{k}$), the process starts with the encoding of $\text{STR}_{x_{k}}$into a latent representation Kingma & Welling (2013): $z=\mathcal{E}(\text{STR}_{x_{k}})$. Gaussian noise is then added to this representation to create $z_{t}$ at timestep $t$. A denoising autoencoder is subsequently trained to predict the Gaussian noise in the latent representation, aiming to reverse the noise addition. The objective is defined as:
@@
where $\epsilon_{\theta}$ is a U-Net Ronneberger et al. (2015) architecture conditioned on a timestep $t$, a text prompt $a$ (action), the visual input $z_{t}$ and the image condition $\text{STR}_{x_{k}}$. To support the condition on images, we follow Brooks et al. (2023) to add additional input channels to the first convolutional layer, concatenating the image condition $\mathcal{E}(\text{STR}_{x_{k}})$ with the noised latent $z_{t}$. After training, our STR predictor is capable of synthesising the next STR ($\text{STR}_{x^{a}_{k+1}}$) for given $\text{STR}_{x_{k}}$ with action instruction $a$.
@@ 3.2.3 GUI-Text Predictor
Given a STR representation generated by our STR predictor, we design a GUI-text predictor to generate plausible text for the text symbols in the STR based on its graphics. Specifically, we first locate the text symbols in the STR by colour matching and boundary detection. This outputs the location of text symbols and their unique ID token, denoted as $\mathcal{T}$. Then we leverage the image processing and task understanding ability of LLM to predict the text content based on its context in STR. This process can be formulated as:
@@
where $\text{STR}_{x^{a}_{k+1}}$ denotes the STR representation of $x^{a}_{k+1}$. $\text{text}_{x^{a}_{k+1}}$ contains the predicted text content for each text symbol, associated with its ID token. This design ensures flexible and accurate text generation tailored to the predicted GUI STRs as the context. Finally, we overlay each text content ($\text{text}_{x^{a}_{k+1}}$) to $\text{STR}_{x^{a}_{k+1}}$ to reconstruct the predicted GUI image ($x^{a}_{k+1}$). To be specific, text symbols are replaced with the corresponding text based on coordinates, with dynamic styling determined by the symbol’s size and surrounding colours. More details are in the Appendix.
@@ 3.3 ViMo Enhanced App Agent
Motivated by that App agents usually face limitations in long-horizon planning to make optimal decisions on action selection Chae et al. (2024), we leverage the proposed ViMo to enhance the decision-making of App agents.
@@
where $a_{se}$ denotes the selected action, and $S(\cdot)$ is the selection model. This procedure is outlined in Algorithm 1. By predicting the next GUI, ViMo provides the agent with the potential outcome of an action, enabling it to make more informed decisions.

### v2

# A Generative Visual GUI World Model for App Agents

@@ Abstract
App agents, which autonomously operate mobile Apps through Graphical User Interfaces (GUIs), have gained significant interest in real-world applications. Yet, they often struggle with long-horizon planning, failing to find the optimal actions for complex tasks with longer steps. To address this, world models are used to predict the next GUI observation based on user actions, enabling more effective agent planning. However, existing world models primarily focus on generating only textual descriptions, lacking essential visual details. To fill this gap, we propose ViMo, the first Visual world Model designed to generate future App observations as images. For the challenge of generating text in image patches, where even minor pixel errors can distort readability, we decompose GUI generation into graphic and text content generation. We propose a novel data representation, the Symbolic Text Representation (STR), to overlay text content with symbolic placeholders while preserving graphics. With this design, ViMo employs a STR Predictor to predict future GUIs’ graphics and a GUI-text Predictor for generating the corresponding text. Moreover, we deploy ViMo to enhance agent-focused tasks by predicting the outcome of different action options. Experiments show ViMo’s ability to generate visually plausible and functionally effective GUIs that enable App agents to make more informed decisions. Project link: https://ai-agents-2030.github.io/ViMo/
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs)  [^1] By LLMs, we refer to the concept of foundation models that accept various input modalities (e.g., visual language models (VLMs), multimodal LLMs (MLLMs)) while producing textual sequences [1]. have unlocked new possibilities for deploying AI agents across diverse fields [2, 3, 4]. A notable application is the smartphone application (App) agents [5, 6], designed to directly interact with Graphical User Interfaces (GUIs) to perform tasks autonomously and efficiently in a mobile operating system.
However, existing agents struggle with making decisions for tasks requiring longer steps [7]. To address this "long-horizon" limitation, an increasing number of studies have introduced world models, which predict how GUIs evolve in response to user actions [8]. Yet, these models typically rely on language to describe future observations. These language-based descriptions often fail to capture the intricate visual details, such as the location and colour of GUI elements, necessary for a precise representation [7]. There is a need for a visual GUI world model capable of comprehensively representing observations in visual modality. Addressing this gap serves as the primary motivation for this work.
Figure 1: GUIs generated by image-based methods (UI-Diffuser [9], TextDiffuser-2 [10], and IP2P [11] fine-tuned on GUI dataset, denoted as IP2P*). Images are cropped for space efficiency.
To build a visual GUI world model capable of generating plausible future GUI observations that are visually consistent with user actions, a straightforward approach involves generating each pixel of a GUI using image generations  [11, 12]. Although these methods demonstrate promising results, such as the GUI graphic generation on the location, style, and colour of GUI elements [9], or scene-text generation in a style that aligns the visual context [10, 13], they still display distortions in the text rendering, particularly for small-sized text where each pixel is critical for accurately identifying and representing the text (see Fig. 1 for an illustration).
@@
We evaluated ViMo in three distinct scenarios to comprehensively demonstrate its effectiveness. First, we assessed its world model capability, where the quality of the generated GUIs was measured using visual similarity, instructional accuracy, and action readiness scores. Each score was examined through both automatic metrics and user studies. These assessments provided a robust and holistic evaluation of how visually precise and contextually plausible the generated GUIs were. Second, we tested ViMo in an agent-focused task to evaluate its benefits for existing App agents and its superiority over other world models. In this setup, given a goal and the current App observation, the agent selected optimal actions to achieve the goal [6]. By accurately predicting the next GUI based on the current observation and an action, ViMo enabled the agent to better anticipate action outcomes and make more informed decisions. This experiment demonstrated the model’s effectiveness in enhancing decision-making for App agents. Finally, we evaluated ViMo’s real-world applicability under two settings: online navigation and zero-shot generalisation. These scenarios assessed the model’s ability to perform in real-time interactions and to generalise to previously unseen Apps, further demonstrating its generalisation capabilities and practical value in dynamic environments.
@@
$\bullet$ Extensive experiments demonstrated the effectiveness of ViMo in both world model evaluation and agent-focused tasks. Specifically, ViMo achieved an average 29.14% and 182.74% relative improvement over existing world models in terms of automatic metrics and user studies, respectively. Moreover, ViMo boosted the step-wise action prediction accuracy of App agents, achieving a 14.07% relative performance gain. In the online navigation setup, ViMo increased the task completion rate from 33.19% to 40.95%, yielding a substantial improvement of 7.76%.
@@ 2.1 App Agent
App agents, powered by LLMs, have made significant strides in automating tasks on mobile Apps [14, 15, 16, 17, 18, 19]. These agents interact with GUIs by emulating human actions such as tapping, typing, and swiping. This enables them to accomplish various user goals, such as searching for products or setting alarms. Approaches in this domain are broadly divided into language-based and multi-modality-based methods. Language-based methods rely on textual description of the App observation and the user goal to generate appropriate actions [5, 20, 21], while multi-modality-based methods enhance this capability by incorporating GUIs for a more comprehensive understanding of the interface [5, 6, 22, 23]. Despite the promising results, these approaches struggle with long-horizon tasks that require multiple interdependent actions and a deep understanding of dynamic environments [7]. For this challenge, recent works [24, 25] introduce the use of real-world emulators to simulate GUI changes resulting from agent actions, enabling App agents to better navigate complex, multi-step scenarios and improve decision-making accuracy. However, the use of real-world emulators faces significant drawbacks, including the high computational costs of setting up emulators for specific mobile systems and the safety risks involved in executing real-world interactions, such as repeatedly sending messages or making purchases. To overcome these, world models have attracted increasing attention as a more efficient alternative [7, 8].
@@ 2.2 World Model
By observing the real world, world models can predict how the environment evolves in response to an action [26, 27]. For instance, video generation models such as Sora [28] simulate future observations by predicting videos given a video observation and GameNGen [29] predicts how a game system will respond to user actions. Notably, the ability to anticipate potential outcomes of actions has proven to be highly beneficial in informing decision-making processes [30, 31, 32, 33]. Inspired by their success, world models have emerged to predict the next observation on websites. These models [7, 8, 34] typically take a website observation and an action as inputs to generate a textual description of the next observation. While websites provide multiple sources of information, including the actual site and their CSS or HTML source files, mobiles present a more limited context, as only the GUIs are typically accessible. Moreover, such text-only descriptions often lack the precise visual details required for accurately predicting future observations, highlighting the need for a visual world model capable of generating high-fidelity future GUI images.
@@ 2.3 GUI Generation
With the rapid advancements in image generation techniques, such as diffusion models [12, 35, 36], generating realistic images has become increasingly feasible. Building on these developments, previous methods have explored generating GUI directly in pixel space. For instance, layout generation methods generate the location of GUI elements [37, 38, 39, 40], scene-text generation methods generate text that aligns with the visual context [10, 13, 41, 42], UI-diffuser [9] fine-tune a stable diffusion model to generate mobile GUIs conditioned on text prompts. For the next GUI generation conditioned on current GUI observation and a user action, it seems straightforward to resort to an image-and-text-conditioned approach [11]. However, we find that pixel-based image generation struggles with rendering text accurately, as even minor pixel prediction errors can lead to distortions, particularly for small-sized text (see Fig. 1 for examples).
@@ 3.2 ViMo: Generative Visual GUI World Model
To tackle the limitation of existing methods [9, 10, 11] in generating visually plausible text for a GUI, as shown in Fig. 1, we propose ViMo, a novel generative visual GUI world model that decouples the graphic and text content generation. As shown in Fig. 2, we first detect and remove all the text in the GUI by overlaying it with a text symbol to create the Symbolic Text Representation (STR). Then a STR predictor is leveraged for determining the STR representation of the next GUI with a pixel-based diffusion process. Finally, a GUI-text predictor is proposed to generate the text content for each symbol using an LLM, followed by a handcrafted design to overplay the text into the STR image to create the next GUI. Their details are specified in the following.
@@ 3.2.1 STR: Symbolic Text Representation
To develop a GUI prediction model that eliminates the need to generate specific text content, we propose the Symbolic Text Representation (STR), where all the text content (actual words) within the GUI image is symbolised (overlayed) with uniform text symbols (placeholders). To be specific, we create an STR representation from a given GUI image with three steps: 1) using an OCR model [43, 44] to detect text within the GUI; 2) masking the detected text by overlaying it with a box filled with white and bordered in black; 3) we leverage an LLM to filter out static text displayed on static GUI elements and preserve it in the image, as it does not involve any semantic evolution or dynamic changes and remains unchanged as part of specific elements such as a keyboard or a clock face. Additionally, we empirically find that predicting this static text with complex spatial patterns poses significant challenges for the LLM.
@@ 3.2.2 STR Predictor
Building on the powerful generative capability of diffusion-based models [12], we introduce a STR predictor specifically trained to understand a given STR and a user action, enabling it to generate the corresponding next STR effectively. In particular, we fine-tune a pre-trained stable diffusion model [12] to predict the next STR, conditioned on the STR of the current GUI and the user action. Given a STR representation ($\text{STR}_{x_{k}}$) extracted from GUI ($x_{k}$), the process starts with the encoding of $\text{STR}_{x_{k}}$into a latent representation [45]: $z=\mathcal{E}(\text{STR}_{x_{k}})$. Gaussian noise is then added to this representation to create $z_{t}$ at timestep $t$. A denoising autoencoder is subsequently trained to predict the Gaussian noise in the latent representation, aiming to reverse the noise addition. The objective is defined as:
@@
where $\epsilon_{\theta}$ is a U-Net [46] architecture conditioned on a timestep $t$, a text prompt $a$ (action), the visual input $z_{t}$ and the image condition $\text{STR}_{x_{k}}$. To support the condition on images, we follow IP2P [11] to add additional input channels to the first convolutional layer, concatenating the image condition $\mathcal{E}(\text{STR}_{x_{k}})$ with the noised latent $z_{t}$. After training, our STR predictor is capable of synthesising the next STR ($\text{STR}_{x^{a}_{k+1}}$) for given $\text{STR}_{x_{k}}$ with action instruction $a$.
@@ 3.2.3 GUI-Text Predictor
Given a STR representation generated by our STR predictor, we design a GUI-text predictor to generate plausible text for the text symbols in the STR based on its graphics. Specifically, we first locate the text symbols in the STR by colour matching and boundary detection. This outputs the location of text symbols, along with their unique ID tokens assigned via enumeration, denoted as $\mathcal{T}$. Then we leverage the image processing and task understanding ability of LLM to predict the text content based on its context in STR. This process can be formulated as:
@@
where $\text{STR}_{x^{a}_{k+1}}$ denotes the STR representation of $x^{a}_{k+1}$. $\text{text}_{x^{a}_{k+1}}$ contains the predicted text content for each text symbol, associated with its ID token. This design ensures flexible and accurate text generation tailored to the predicted GUI STRs as the context. Finally, we overlay each text content ($\text{text}_{x^{a}_{k+1}}$) to $\text{STR}_{x^{a}_{k+1}}$ to reconstruct the predicted GUI image ($x^{a}_{k+1}$). To be specific, text symbols are replaced with the corresponding text based on coordinates, with dynamic styling determined by the symbol’s size and surrounding colours. More details are provided in the Appendix.
@@ 3.3 ViMo Enhanced App Agent
Motivated by that App agents usually face limitations in long-horizon planning to make optimal decisions on action selection [7], we leverage the proposed ViMo to enhance the decision-making of App agents.
@@
where $a_{se}$ denotes the selected action, and $S(\cdot)$ is the selection model. This procedure is outlined in Algorithm 1. By predicting the next GUI, ViMo provides the agent with the potential outcome of an action, enabling it to make more informed decisions. We build the selection model to identify the best action in two steps. First, we query an LLM to evaluate all the action candidates, providing a judgment—either valid or invalid—and a confidence score for each action. Second, we query the LLM again to select the best action from the two highest-scoring actions. This process is motivated by our observation that, in over $70\%$ of tasks, the difference between the top two scores is equal to or less than $0.1$, indicating that both are likely optimal. By explicitly prompting the LLM to compare the top candidates, we go beyond coarse scoring and enable more detailed decision-making.
Table 1: GUI quality evaluation. $s_{gc}$ indicates the GUI consistency, $s_{ia}$ instructional accuracy and $s_{ar}$ action readiness score. $s_{h}$ indicates the harmonic average between the 3 metrics. $\Delta s_{h}$ is the relative performance gains of our ViMo over other methods. IP2P* denotes finetuing of IP2P on our dataset.
Method Automatic Metric User Study
$s_{gc}$ $s_{ia}$ $s_{ar}$ $s_{h}$ $\Delta s_{h}$ $s_{gc}$ $s_{ia}$ $s_{ar}$ $s_{h}$ $\Delta s_{h}$
HTML-vision 0.70 85.77 62.79 0.72 5.39% 0.31 11.32 9.01 0.23 282.61%
IP2P* 0.74 63.57 70.15 0.69 10.20% 0.82 58.92 52.81 0.63 39.68%
UI-diffuser 0.60 39.61 38.75 0.44 71.82% 0.36 14.32 8.56 0.27 225.93%
ViMo (Ours) 0.74 75.39 78.68 0.76 - 0.89 91.12 84.71 0.88 -